### v1

# Quantum Brownian Motion as a Classical Stochastic Process in Phase Space

Abstract
We establish that the exact quantum dynamics of a Brownian particle in the Caldeira-Leggett model can be mapped, at any temperature, onto a classical, non-Markovian stochastic process in phase space. Starting from a correlated thermal equilibrium state between the particle and bath, we prove that this correspondence is exact for quadratic potentials under arbitrary quantum state preparations of the particle itself. For more general, smooth potentials, we identify and exploit a natural small parameter: the density matrix becomes strongly quasidiagonal in the coordinate representation, with its off-diagonal width shrinking as the bath’s spectral cutoff increases, providing a controlled parameter for accurate approximation. The framework is fully general: arbitrary initial quantum states—including highly non-classical superpositions—are incorporated via their Wigner functions, which serve as statistical weights for trajectory ensembles. Furthermore, the formalism naturally accommodates external manipulations and measurements modeled by preparation functions acting at arbitrary times, enabling the simulation of complex driven-dissipative quantum protocols.
I Introduction
Brownian motion—the erratic dynamics of a microscopic system interacting with a vast environment—is a cornerstone of statistical physics. Its classical description in terms of a Langevin equation Langevin1908 , combining friction and random noise, is ubiquitous across chemical kinetics, biophysics, and materials science Hanggi1990 . On smaller scales where quantum effects become significant, such as in low-temperature defect dynamics, quantum diffusion or tunneling Leggett1984QuantumTunneling ; BEC ; torres2022surface ; gottwald2015applicability ; Caldeira1983 , a fundamental question arises: how can Brownian motion be consistently quantized? This problem, known as quantum Brownian motion (QBM), is notoriously difficult. The principles of quantum mechanics forbid a naive quantization of classical dissipative equations, leaving a microscopic description of the system coupled to a macroscopic environment as the only rigorous path. However, the exponential growth of the Hilbert space dimension makes direct simulation intractable, creating a pressing need for a reduced description in terms of the Brownian particle alone.
The first microscopically consistent model for QBM was introduced by Caldeira and Leggett (CL) Caldeira1983path , representing the environment as a bath of independent harmonic oscillators linearly coupled to the particle. A foundational and well-known result of this model is that the Heisenberg equations of motion for the particle operators are formally identical to the classical generalized Langevin equation Weiss_2012 ; gardiner2004quantum . This formal correspondence suggests that the quantum dynamics might be representable in classical terms, yet translating this operator identity into a practical, all-temperature description of the quantum state has remained a challenge.
Early attempts to derive reduced equations of motion often resorted to approximations valid only in limiting regimes. For instance, master equations of Lindblad form Lindblad1976 , which are common in quantum optics, can be derived for QBM but are strictly valid only in the high-temperature limit ($k_{B}T\gg\hbar\Lambda$, where $\Lambda$ is a bath cutoff) breuer . At low temperatures, these equations fail to capture essential quantum features such as zero-point fluctuations and can even violate physical principles Vacchini2000 ; Vacchini2001 . Despite this limitation, their relative simplicity has sustained their use in the literature. The exact path-integral solutions Feynman1965 ; Feynman1963 for the CL model, while powerful Grabert1988 ; Schramm1986 ; Hu1992 , are analytically cumbersome and limited to quadratic potentials, hindering their application to realistic anharmonic systems. This motivates the search for a more direct computational framework that leverages the intrinsic structure of the CL model.
In this work, we demonstrate that the exact quantum dynamics of the CL model admits a complete and practical representation as a classical, non-Markovian stochastic process in phase space, valid at all temperatures. We prove that, starting from a correlated thermal equilibrium state, the reduced quantum dynamics for a particle in a quadratic potential is identical to an ensemble of trajectories governed by a classical generalized Langevin equation. The quantum nature of the environment is encoded exactly in the statistics of a colored noise that obeys the quantum fluctuation-dissipation theorem. For smooth, non-quadratic potentials, we identify a natural small parameter—the off-diagonal width (coherence length) of the equilibrium reduced density matrix—which shrinks with increasing bath spectral cutoff and controls the accuracy of the approximation.
The primary value of this mapping is its profound simplification: it reduces the formidable problem of simulating an open quantum system with infinitely many degrees of freedom to the tractable task of generating and averaging classical stochastic trajectories. Our associated numerical technique accommodates arbitrary initial quantum states—including non-classical superpositions—via their Wigner representations, and can incorporate external manipulations through preparation functions. This offers a conceptually straightforward and versatile alternative to more complex contemporary methods, providing a unified computational tool that works seamlessly from the classical high-temperature regime down to the deeply quantum, zero-temperature limit.
The remainder of this paper is structured as follows: In Sec. II, we review the Caldeira-Leggett model and the exact path-integral framework. In Sec. III, we present the core result—the exact classical-stochastic mapping—and detail its proof. Section IV describes the numerical implementation. In Sec. V, we validate the method against known analytical results and contrast it with high-temperature master equations, highlighting the crucial all-temperature behavior. Section VI discusses the extension to non-quadratic potentials via the identified small parameter. We conclude in Sec. VII.
II Microscopic model of quantum Brownian motion
II.1 Classical Caldeira-Leggett Model
The paradigmatic model for quantum Brownian motion is the Caldeira-Leggett (CL) model Caldeira1983path , which represents a one-dimensional Brownian particle of mass $m$ bilinearly coupled to a macroscopic environment. The environment is modeled as a bath of independent harmonic oscillators. The Hamiltonian for the total system is
$\begin{gathered}H=\frac{p^{2}}{2m}+V(x)+\sum_{i}\left[\frac{p_{i}^{2}}{2m_{i}}+\frac{1}{2}m_{i}\omega_{i}^{2}x_{i}^{2}\right]\\ -x\sum_{i}c_{i}x_{i}+x^{2}\sum_{i}\frac{c_{i}^{2}}{2m_{i}\omega_{i}^{2}},\end{gathered}$ (1)
where $x,p$ are the coordinate and momentum of the particle, and $x_{i},p_{i},m_{i},\omega_{i}$ are the coordinates, momenta, masses, and frequencies of the bath oscillators. The constants $c_{i}$ quantify the coupling strength. The first two terms constitute the Hamiltonian of the isolated particle (we consider free Brownian motion with $V(x)=0$). The third term is the Hamiltonian of the isolated bath. The fourth term describes the linear system-bath interaction, and the final counter-term is necessary to ensure that there are no additional forces acting on the free particle at rest Caldeira1983 ; Weiss_2012 .
II.1.1 Reduced Equation of Motion
The classical equations of motion derived from Hamilton’s equations are
$m\ddot{x}(t)=\sum_{i}c_{i}x_{i}(t)-x(t)\sum_{i}\frac{c_{i}^{2}}{m_{i}\omega_{i}^{2}},$ (2)
$m_{i}\ddot{x}_{i}(t)+m_{i}\omega_{i}^{2}x_{i}(t)=c_{i}x(t).$ (3)
Equation (3) for the bath oscillators is a linear inhomogeneous equation. Its formal solution, substituting the particle’s trajectory $x(t)$, is
$x_{i}(t)=x_{i}(0)\cos(\omega_{i}t)+\frac{p_{i}(0)}{m_{i}\omega_{i}}\sin(\omega_{i}t)$ (4)
$+\frac{c_{i}}{m_{i}\omega_{i}}\int_{0}^{t}\sin[\omega_{i}(t-\tau)]x(\tau)\,d\tau.$
Substituting Eq. (4) into Eq. (2) yields a closed, non-Markovian equation for the particle alone:
$m\ddot{x}(t)=-\int_{0}^{t}M(t-\tau)\dot{x}(\tau)\,d\tau-x(0)M(t)+\xi(t).$ (5)
This is a generalized Langevin equation (GLE). Its constituents are:
  • The memory kernel (friction kernel):
    $M(t)=\sum_{i}\frac{c_{i}^{2}}{m_{i}\omega_{i}^{2}}\cos(\omega_{i}t).$ (6)
  • The stochastic force (noise):
    $\xi(t)=\sum_{i}c_{i}\left[x_{i}(0)\cos(\omega_{i}t)+\frac{p_{i}(0)}{m_{i}\omega_{i}}\sin(\omega_{i}t)\right].$ (7)
The term x(0)M(t)-x(0)M(t) arises from the non-translational-invariant PhysRev2 preparation of the initial bath state relative to the particle. It is non-physical for describing steady-state Brownian motion and is typically removed by assuming the bath was in equilibrium with the particle held at position $x(0)$ for all past times. This is equivalent to redefining the initial bath coordinates as $x_{i}(0)\to x_{i}(0)+\frac{c_{i}}{m_{i}\omega_{i}^{2}}x(0)$, which cancels the term exactly. We adopt this physically consistent preparation, yielding the standard GLE:
$m\ddot{x}(t)=-\int_{0}^{t}M(t-\tau)\dot{x}(\tau)\,d\tau+\xi(t).$ (8)
II.1.2 Spectral Density and Thermodynamic Limit
In the thermodynamic limit of a continuous bath, the system properties are encoded in the spectral density
$J(\omega)=\frac{\pi}{2}\sum_{i}\frac{c_{i}^{2}}{m_{i}\omega_{i}}\delta(\omega-\omega_{i}).$ (9)
The memory kernel is then expressed as
$M(t)=\frac{2}{\pi}\int_{0}^{\infty}d\omega\,\frac{J(\omega)}{\omega}\cos(\omega t).$ (10)
For an ohmic bath with an exponential cutoff on $\Lambda$ (corresponding to a time scale $\varepsilon=\Lambda^{-1}$), $J(\omega)=\gamma m\omega e^{-\varepsilon\omega}$, where $\gamma$ is the friction coefficient. The corresponding memory kernel is $M(t)=(2m\gamma/\pi)\varepsilon/(\varepsilon^{2}+t^{2})$. In the Markovian limit $\gamma\varepsilon\to 0$, $M(t)\to 2m\gamma\delta(t)$, and Eq. (8) reduces to the standard Langevin equation:
$m\ddot{x}(t)=-m\gamma\dot{x}(t)+\xi(t).$ (11)
II.1.3 Fluctuation-Dissipation Theorem
The stochastic force $\xi(t)$ derives from the uncertain initial conditions $\{x_{i}(0),p_{i}(0)\}$ of the bath. Assuming these are sampled from a Gibbs distribution at temperature $T$, i.e., $\rho_{B}\propto\exp(-H_{B}/k_{B}T)$, the noise is Gaussian with zero mean and its autocorrelation function is given by the classical fluctuation-dissipation theorem (FDT):
$\langle\xi(t)\xi(s)\rangle=k_{B}T\,M(t-s).$ (12)
In the Markovian limit, this becomes $\langle\xi(t)\xi(s)\rangle=2m\gamma k_{B}T\,\delta(t-s)$.
II.2 Quantum Caldeira-Leggett Model
The quantization of the CL model proceeds by promoting the classical phase-space variables to operators satisfying canonical commutation relations: $x,p]=i\hbar$, $x_{i},p_{j}]=i\hbar\delta_{ij}$, with all other commutators vanishing.
II.2.1 Joint State and Preparation
The total system’s quantum state is described by a density matrix $\rho_{SB}$. We consider the physically relevant scenario where, at an initial time $t=0$, the total system is prepared starting from a joint thermal equilibrium state at inverse temperature $\beta=1/(k_{B}T)$:
$\rho_{SB}^{\beta}=Z^{-1}e^{-\beta H},$ (13)
where $H$ is the full Hamiltonian in Eq. (1) and $Z$ is the partition function. This state contains all system-bath correlations consistent with equilibrium. An arbitrary preparation of the Brownian particle’s state is then enacted via an operation acting solely on its Hilbert space Grabert1988 . This is described by a set of preparation operators $\{O_{j}\}$, yielding the initial state
$\rho_{SB}(0)=\sum_{j}O_{j}\,\rho_{SB}^{\beta}\,O_{j}^{\dagger}.$ (14)
This formalism can be generalized to describe interventions (e.g., measurements or controlled manipulations) at intermediate times $t_{k}$.
II.2.2 Reduced Quantum Dynamics with Intermediate Interventions
The formalism of preparation operators naturally extends to describe a sequence of interventions—such as measurements, unitary kicks, or state preparations—applied to the particle at specific intermediate times. Consider a sequence of times $0<t_{1}<t_{2}<\dots<t_{N}$. At each time $t_{k}$, an intervention is described by a set of operators $\{O_{j}^{(k)}\}$ acting solely on the particle’s Hilbert space. The quantum state evolves as follows:
  1. 1.
    Initial Preparation: Starting from the joint thermal state $\rho_{SB}^{\beta}$ from Eq. (13), an initial preparation is performed at $t=0$:
    $\rho_{SB}(0^{+})=\sum_{j_{0}}O_{j_{0}}^{(0)}\,\rho_{SB}^{\beta}\,\big(O_{j_{0}}^{(0)}\big)^{\dagger}.$ (15)
  2. 2.
    Unitary Evolution: The state evolves unitarily under the full Hamiltonian until the first intervention:
    $\rho_{SB}(t_{1}^{-})=U(t_{1})\,\rho_{SB}(0^{+})\,U^{\dagger}(t_{1}),\quad\text{where }U(t)=e^{-iHt/\hbar}.$ (16)
  3. 3.
    Intervention at $t_{k}$: At time $t_{k}$, the intervention operators are applied:
    $\rho_{SB}(t_{k}^{+})=\sum_{j_{k}}O_{j_{k}}^{(k)}\,\rho_{SB}(t_{k}^{-})\,\big(O_{j_{k}}^{(k)}\big)^{\dagger}.$ (17)
    The state then continues its unitary evolution until the next intervention or the final time.
  4. 4.
    Final State and Observables: After the last intervention at $t_{N}$, the state evolves to the final time $t_{f}$:
    $\rho_{SB}(t_{f})=U(t_{f}-t_{N})\,\rho_{SB}(t_{N}^{+})\,U^{\dagger}(t_{f}-t_{N}).$ (18)
    The physically relevant object is the reduced density matrix of the particle at $t_{f}$, obtained by tracing out the bath:
[…]
    The expectation value of any particle observable $\hat{A}$ at an arbitrary time $t$ within the interval $0,t_{f}]$ is defined by the state at that time. The observable $\hat{A}$ acts solely on the particle’s Hilbert space. The time $t$ may precede, follow, or lie between the intervention times $\{t_{k}\}$. The general expression is
[…]
This framework provides a complete prescription for computing quantum observables under arbitrary sequences of preparations and interventions within the CL model. The key result of our work is that this complex, non-Markovian quantum evolution can be mapped exactly onto a classical stochastic process for quadratic potentials, as detailed in the following sections.
III Wigner Picture and the Classical-Quantum Correspondence
The full Hamiltonian of the Caldeira-Leggett model is quadratic in both the particle and bath coordinates. This fundamental property leads to a profound simplification in the phase-space description of the quantum dynamics: the quantum Liouville equation for the Wigner function of the total system is formally identical to the classical Liouville equation for the joint probability distribution in phase space Weiss_2012 .
Wigner function is defined by
$W(r,p,t)=\int dqe^{-\frac{i}{\hbar}pq}\rho(r,q,t),$ (21)
where we introduced ”center mass” variables $r=\frac{x+x^{\prime}}{2},\,\,q=x-x^{\prime}$ for density matrix $\rho(x,x^{\prime},t)$ in coordinate representation.
III.1 From Wigner Dynamics to Stochastic Trajectories
Let $W^{SB}(W^{SB}(r,p,\{r_{i},p_{i}\},t)$ denote the Wigner function of the total system (particle + bath). $(r,p)$ - particle, $(r_{i},p_{i})$ - all oscillators. For a quadratic Hamiltonian, its evolution is given by
$\frac{\partial W^{SB}}{\partial t}=\{H,W^{SB}\}_{\text{P.B.}},$ (22)
where $\{\cdot,\cdot\}_{\text{P.B.}}$ is the classical Poisson bracket. Equation (22) is the classical Liouville equation. Its solution can be represented by an ensemble of deterministic trajectories in the total phase space, evolving under Hamilton’s equations derived from $H$.
This establishes a rigorous trajectory-based interpretation of the exact quantum dynamics for the CL model. The procedure is as follows:
  1. 1.
    At a distant past time t=Tt=-T (with $T\to\infty$), we assume a factorized, positive initial Wigner function:
    $W^{SB}(-T)=W(r,p;-T)\times\prod_{i}W_{i}^{\beta}(r_{i},p_{i}).$ (23)
    Here, $W$ is an arbitrary (positive) Wigner function for the Brownian particle, and $W_{i}^{\beta}$ is the thermal Wigner function for the $i$-th bath oscillator at inverse temperature $\beta=1/(k_{B}T)$, which for a harmonic oscillator is a positive Gaussian. Crucially, evolving the system from this factorized state at t=Tt=-T to $t=0$ under the full interacting Hamiltonian allows the particle and the bath to fully equilibrate. In the limit $T\to\infty$, the state at $t=0$ converges to the true correlated thermal state $\rho_{SB}^{\beta}$ from Eq. (13), establishing all necessary system-environment correlations.
  2. 2.
    An ensemble of initial conditions $(r^{(j)},p^{(j)},\{r_{i}^{(j)},p_{i}^{(j)}\})$ is sampled from the distribution $W^{SB}(-T)$.
  3. 3.
    For each sample $j$, the classical equations of motion (Hamilton’s equations derived from the full $H$) are solved numerically from t=Tt=-T to the desired final time.
  4. 4.
    Expectation values of system observables at any time $t$ are computed by averaging the corresponding Weyl symbol $O_{W}(r,p)$ over this ensemble of classical trajectories:
[…]
III.2 The Reduced Stochastic Process for the Brownian Particle
The Hamiltonian structure of the CL model allows the bath degrees of freedom to be eliminated analytically from the joint trajectories, exactly as in the classical case (Section 2.1). For each sampled trajectory, the motion of the Brownian particle obeys the generalized Langevin equation (GLE):
$m\ddot{x}(t)=-\int_{-T}^{t}M(t-\tau)\dot{x}(\tau)d\tau+\xi^{(j)}(t),$ (25)
where the memory kernel $M(t)$ is defined in Eq. (10). The crucial difference from the classical case lies in the statistics of the stochastic force $\xi^{(j)}(t)$.
The force $\xi^{(j)}(t)$ is constructed from the sampled initial conditions of the bath oscillators $\{r_{i}^{(j)}(-T),p_{i}^{(j)}(-T)\}$:
$\xi^{(j)}(t)=\sum_{i}c_{i}\Big[$ $r_{i}^{(j)}(-T)\cos(\omega_{i}(t+T))$ (26)
$+\frac{p_{i}^{(j)}(-T)}{m_{i}\omega_{i}}\sin(\omega_{i}(t+T))\Big].$
Because these initial conditions are drawn from the quantum thermal Wigner functions $W_{i}^{\beta}$, the ensemble-averaged noise correlation is given by the quantum fluctuation-dissipation theorem:
$\langle\xi(t)\xi(s)\rangle_{\text{ens}}=\frac{1}{\pi}\int_{0}^{\infty}d\omega\,J(\omega)\hbar\omega\\ \times\left[\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\cos[\omega(t-s)]-i\sin[\omega(t-s)]\right].$ (26)
The symmetric (real) part of this correlation, which governs observable averages, is
$\frac{1}{2}\langle\{\xi(t),\xi(s)\}\rangle_{\text{ens}}=\frac{1}{\pi}\int_{0}^{\infty}d\omega\,J(\omega)\hbar\omega\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\cos[\omega(t-s)].$ (27)
At $T=0$, $\coth(\hbar\omega/(2k_{B}T))\to 1$, leading to zero-point fluctuation noise. For an ohmic bath $J(\omega)=\gamma m\omega e^{-\varepsilon\omega}$, this yields the non-Markovian correlation function
$\langle\xi(t)\xi(0)\rangle_{T=0}=\frac{m\gamma\hbar}{\pi}\frac{\varepsilon^{2}-t^{2}}{(\varepsilon^{2}+t^{2})^{2}},$ (28)
which is clearly distinct from the classical white-noise limit.
Thus, the exact quantum dynamics of the particle is reproduced by an ensemble of trajectories solving the classical GLE (25), where the only ”quantum” ingredient is the noise statistics derived from the initial quantum Wigner distribution of the bath.
III.3 Incorporating State Preparation and Interventions
The preparation formalism of Section 2.2 is incorporated naturally within this trajectory picture. A preparation or intervention at time $t_{k}$, described by operators $\{O_{\alpha}^{(k)}\}$, corresponds to a non-classical update of the particle’s Wigner function. In our stochastic framework, this is implemented by assigning a complex weight $w_{\alpha}^{(j)}$ to each trajectory $j$ at time $t_{k}$.
The weight is derived from the preparation function $\lambda$ in the Wigner representation. For a trajectory that has phase-space coordinates $(\bar{r},\bar{p})$ just before $t_{k}$ and $(r_{0},p_{0})$ just after, the weight contributed by the intervention is
$w^{(j)}(t_{k})=\lambda(r_{0},p_{0}|\bar{r},\bar{p}).$ (29)
For a sequence of interventions, the total weight for a trajectory is the product of weights from all events. The expectation value of an observable at time $t$ is then computed by a weighted average over the ensemble:
$\langle\hat{O}(t)\rangle=\frac{\sum_{j=1}^{N}\left(\prod_{\{t_{k}<t\}}w^{(j)}(t_{k})\right)O_{W}\left(r^{(j)}(t),p_{S}^{(j)}(t)\right)}{\sum_{j=1}^{N}\left(\prod_{\{t_{k}<t\}}w^{(j)}(t_{k})\right)}.$ (30)
This formulation provides a complete and practical method for simulating the non-Markovian quantum dynamics of the Brownian particle under arbitrary sequences of preparations and measurements, provided the potential is quadratic. The numerical implementation of this stochastic approach is detailed in the next section.
IV Numerical Method: Stochastic Monte Carlo Simulation
The theoretical correspondence established in the previous sections—mapping the quantum dynamics of the Caldeira-Leggett model to a classical stochastic process—lends itself directly to numerical implementation via a Monte Carlo technique. The core algorithm involves: (i) sampling initial conditions from the appropriate equilibrium distributions, (ii) generating stochastic trajectories by solving the generalized Langevin equation (GLE), and (iii) computing quantum observables via weighted ensemble averages that account for state preparations.
IV.1 Generation of Quantum Noise
The key ingredient for simulating the quantum GLE is the generation of stochastic force trajectories $\xi^{(j)}(t)$ whose two-time correlation satisfies the quantum fluctuation-dissipation theorem:
$\langle\xi(t)\xi(s)\rangle=\frac{1}{\pi}\int_{0}^{\infty}d\omega\,J(\omega)\hbar\omega\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\cos[\omega(t-s)].$ (31)
A numerically efficient method is to synthesize $\xi(t)$ in the frequency domain. We express the noise as
$\xi(t)=\sqrt{\frac{\hbar}{2\pi}}\int_{-\infty}^{\infty}d\omega\,\sqrt{J(|\omega|)S_{\text{th}}(|\omega|)}\,z(\omega)e^{-i\omega t},$ (32)
where $S_{\text{th}}(\omega)=\coth(\hbar\omega/(2k_{B}T))$ and $z(\omega)$ is a complex Gaussian auxiliary noise with statistics
$z^{*}(\omega)=z(-\omega),\quad\langle z(\omega)z^{*}(\omega^{\prime})\rangle=\delta(\omega-\omega^{\prime}),\quad\langle z(\omega)z(\omega^{\prime})\rangle=0.$ (33)
The discretized numerical implementation uses a finite frequency grid $\omega_{k}=k\Delta\omega$ for $k=-N,\dots,N$. For each $k$, we generate independent complex random variables $\eta_{k}$ and $\zeta_{k}$ from a standard normal distribution $\mathcal{N}(0,1)$ and construct
$z_{k}=\frac{1}{\sqrt{2}}(\eta_{k}+i\zeta_{k}),\quad\text{for }k>0,$ (34)
with $z_{0}$ being real ($\eta_{0}$) and $z_{-k}=z_{k}^{*}$ to satisfy Eq. (33). The noise time series at discrete times $t_{n}$ is then computed via the inverse discrete Fourier transform:
$\xi(t_{n})=\sqrt{\frac{\hbar\Delta\omega}{2\pi}}\sum_{k=-N}^{N}\sqrt{J(|\omega_{k}|)S_{\text{th}}(|\omega_{k}|)}\,z_{k}\,e^{-i\omega_{k}t_{n}}.$ (35)
This procedure generates a stationary Gaussian process whose ensemble average reproduces Eq. (31) with accuracy controlled by the spectral resolution $\Delta\omega$ and cutoff $N\Delta\omega$.
IV.2 Trajectory Generation and Equilibrium Initialization
For a given noise realization $\xi^{(j)}(t)$, the particle’s trajectory $(x^{(j)}(t),p^{(j)}(t))$ is obtained by numerically integrating the classical GLE (derived from Eq. (25) for a free particle or in a potential $V(x)$):
$m\ddot{x}^{(j)}(t)=-\frac{\partial V}{\partial x}\bigg|_{x^{(j)}(t)}-\int_{t_{0}}^{t}d\tau\,M(t-\tau)\dot{x}^{(j)}(\tau)+\xi^{(j)}(t).$ (36)
To ensure the initial state at the simulation start time $t=0$ corresponds to the correlated thermal equilibrium $\rho_{SB}^{\beta}$, we employ the following protocol:
  1. 1.
    Begin simulations at a sufficiently distant past time $t_{\text{start}}=-T_{\text{eq}}$, with $T_{\text{eq}}$ chosen to be several times longer than the bath’s memory time $\tau_{M}$ and the system’s relaxation time $\gamma^{-1}$.
  2. 2.
    Initialize the particle with arbitrary coordinates and momentum (e.g., $x(-T_{\text{eq}})=0,p(-T_{\text{eq}})=0$).
  3. 3.
    Evolve the system under Eq. (36) from $t_{\text{start}}$ to $t=0$. This equilibration period allows the system-bath correlations to fully develop. The state at $t=0$ is then statistically indistinguishable from the true thermal equilibrium state $\rho_{SB}^{\beta}$.
  4. 4.
    Store the particle’s phase-space coordinates $(x^{(j)}(0^{-}),p^{(j)}(0^{-}))$ at the end of this equilibration run.
This initialization procedure implicitly samples from the correct Wigner distribution $W^{SB}(r,p,\{r_{i},p_{i}\},0)$ corresponding to $\rho_{SB}^{\beta}$, circumventing the need for its explicit construction.
IV.3 Monte Carlo Sampling of Preparations and Weighted Averages
The action of a preparation operator at $t=0$, described by the preparation function $\lambda(r_{i},p_{i}|\bar{r},\bar{p})$ in the Wigner representation, is implemented as a reweighting and resampling step on the ensemble of equilibrated trajectories.
  1. 1.
    For each trajectory $j$, the coordinates $(\bar{r}^{(j)},\bar{p}^{(j)})=(x^{(j)}(0^{-}),p^{(j)}(0^{-}))$ serve as the input to the preparation function.
  2. 2.
    A new set of post-preparation coordinates $(r_{i}^{(j)},p_{i}^{(j)})$ is sampled. In this work, we employ a simple rejection sampling method: candidate points $(r,p)$ are drawn uniformly from a region encompassing the essential support of the target distribution (e.g., within several standard deviations of its mean), and accepted with a probability proportional to $|\lambda(r,p|\bar{r}^{(j)},\bar{p}^{(j)})|$.
  3. 3.
    Each trajectory is assigned a complex weight
    $w^{(j)}=\lambda(r_{0}^{(j)},p_{0}^{(j)}|\bar{r}^{(j)},\bar{p}^{(j)}).$ (37)
  4. 4.
    The trajectory is then propagated forward from the new initial condition $(r_{0}^{(j)},p_{0}^{(j)})$ at $t=0^{+}$ under Eq. (36), generating its future evolution.
For a sequence of interventions at times $\{t_{k}\}$, this reweighting/resampling procedure is applied at each intervention time using the corresponding preparation function. The expectation value of a particle observable $\hat{O}$ with Weyl symbol $O_{W}(x,p)$ at any time $t>0$ is computed via the weighted Monte Carlo average over the ensemble of $N_{\text{traj}}$ trajectories:
$\langle\hat{O}(t)\rangle=\frac{\sum_{j=1}^{N}w^{(j)}\,O_{W}\left(x^{(j)}(t),p^{(j)}(t)\right)}{\sum_{j=1}^{N}w^{(j)}}.$ (38)
The statistical error scales as $1/\sqrt{N}$, as is typical for Monte Carlo methods. This numerical framework provides a direct and computationally tractable path for simulating the exact non-Markovian quantum dynamics of Brownian motion for arbitrary temperatures and state preparations.
V Results and Comparison with Established Methods
We now validate our stochastic numerical method and use it to investigate phenomena that are intractable for standard high-temperature approximations. First, we benchmark the method against exact analytical results for Gaussian preparations. We then employ it to study the decoherence of a non-classical state (a Schrödinger cat), comparing the exact quantum dynamics with the predictions of the common high-temperature master equation. Finally, we analyze the distinct thermalization behavior at low temperatures.
V.1 Benchmark: Gaussian State Preparation
As a first test, we consider a Gaussian preparation that localizes the particle’s position Grabert1988 . In coordinate representation, the preparation function is
$\lambda(r_{0},\bar{r},q_{0},\bar{q})=\delta(q_{0}-\bar{q})\delta(r_{0}-\bar{r})\frac{1}{\sqrt{2\pi\sigma_{0}^{2}}}\exp\left(-\frac{r_{0}^{2}}{2\sigma_{0}^{2}}-\frac{q_{0}^{2}}{8\sigma_{0}^{2}}\right),$ (39)
which corresponds to a state centered at $\langle x\rangle=0$ with initial spatial variance $\sigma_{0}^{2}$. The uncertainty principle mandates a corresponding increase in momentum variance, reflected in the Wigner representation:
$\lambda(r_{0},\bar{r},p_{0},\bar{p})=\delta(r_{0}-\bar{r})\frac{1}{2(2\pi)^{2}}\exp\left(-\frac{r_{0}^{2}}{2\sigma_{0}^{2}}-\frac{2\sigma_{0}^{2}}{\hbar^{2}}(p_{0}-\bar{p})^{2}\right).$ (40)
For this preparation within the quadratic CL model, the mean squared displacement has an exact analytical expression Grabert1988 :
$\sigma^{2}(t)=\langle x^{2}(t)\rangle=\sigma_{0}^{2}+d^{2}(t)+\frac{A^{2}(t)}{\sigma_{0}^{2}},$ (41)
where $d^{2}(t)=\langle(x(t)-x(0))^{2}\rangle$ is the mean squared displacement of the thermal particle and $A(t)=(1/2i)\langle[x(t),x(0)]\rangle$ is the commutator response function.
Figure 1 shows the time evolution of the coordinate dispersion $\sigma^{2}(t)$ at zero temperature ($T=0$), computed using our stochastic method (data points) compared to the analytical formula Eq. (41) (solid line). The parameters are $\sigma_{0}=1$, $\gamma=\pi/2$, $\varepsilon=0.5$, and $m=\hbar=1$. The perfect agreement validates the core numerical implementation: our method of sampling noise from the zero-point spectral density and weighting trajectories via the Wigner preparation function accurately reproduces exact quantum results.
Figure 1: Time evolution of the coordinate dispersion $\sigma^{2}(t)$ for a Gaussian preparation at $T=0$. Points: stochastic Monte Carlo simulation. Solid line: analytical result from Eq. (41). The agreement validates the numerical method. Parameters: $\sigma_{0}=1$, $\gamma=\pi/2$, $\varepsilon=0.5$, $m=\hbar=1$.
V.2 Decoherence Dynamics: Schrödinger Cat State
To demonstrate the capability of our method for non-classical states and to highlight the limitations of high-temperature approximations, we study the decoherence of a Schrödinger cat state. The state is a superposition of two spatially separated Gaussian wave packets:
$\begin{gathered}|\psi\rangle=\frac{1}{\sqrt{\mathcal{N}}}(|+\rangle+|-\rangle),\\ \langle x|\pm\rangle=\frac{1}{(2\pi\sigma^{2})^{1/4}}\exp\left(-\frac{(x\pm x_{0})^{2}}{4\sigma^{2}}\right),\end{gathered}$ (42)
where $\mathcal{N}=2(1+\langle-|+\rangle)$. The corresponding preparation function in the Wigner representation factorizes: $\lambda(r,p|\bar{r},\bar{p})=\mathcal{Z}^{-1}W_{\text{Cat}}(r,p)W_{\text{Cat}}(\bar{r},\bar{p})$.
[…]
$O_{W}(x,p)=4\exp\left(-\frac{x^{2}}{2\sigma^{2}}-2\sigma^{2}p^{2}\right)\cos(2x_{0}p).$ (43)
We compare three distinct dynamics in Fig. 2:
  1. 1.
    Exact Quantum Dynamics (Our method): The colored noise correlation uses the full quantum FDT, Eq. (31).
  2. 2.
    High-Temperature White Noise: The noise correlation uses the classical FDT, $\langle\xi(t)\xi(0)\rangle\propto 2m\gamma k_{B}T\delta(t)$, corresponding to the standard Markovian limit.
  3. 3.
    High-Temperature Master Equation (ME): Dynamics from the common Caldeira-Leggett master equation Caldeira1983 ; breuer :
    $\dot{\rho}=-\frac{i}{\hbar}[H_{0},\rho]-\frac{i\gamma}{2\hbar}[x,\{p,\rho\}]-\frac{m\gamma k_{B}T}{\hbar^{2}}[x,[x,\rho]].$ (44)
The parameters are $\gamma=\pi/2$, $\varepsilon=0.01$, and $T=1$ (in units where $\hbar=1$). Crucially, for these parameters, $k_{B}T\approx 1$ is not much greater than the effective cutoff energy $\hbar/\varepsilon\approx 100$, violating the condition ($k_{B}T\gg\hbar/\varepsilon$) required for the Markovian/high-temperature approximation to be valid Ferialdi .
Figure 2 reveals significant differences. The high-temperature ME and white-noise stochastic process predict nearly identical, relatively slow decoherence. In stark contrast, the exact quantum dynamics exhibits much faster initial decoherence. This is a direct manifestation of zero-point and low-temperature fluctuations in the bath, which are neglected in the classical FDT. Our method, which incorporates the full quantum noise spectrum, is essential to capture this correct physics.
Figure 2: Decoherence of a Schrödinger cat state. The observable $\langle O\rangle$ (Eq. (43)) decays fastest for the exact quantum noise (solid line). The high-temperature master equation (dashed) and white-noise process (dotted) give similar, slower decoherence, failing to capture the enhanced quantum fluctuations at low temperature. Parameters: $\gamma=\pi/2$, $\varepsilon=0.01$, $T=1$, $m=\hbar=1$.
V.3 Low-Temperature Thermalization and the Role of the Cutoff
The failure of the high-temperature approximation is further elucidated by examining momentum thermalization. Figure 3 shows the time evolution of $\langle p^{2}(t)\rangle$ starting from an initial value of zero, for the same parameters as in Fig. 2. The exact quantum result (from both analytical integration breuer and our stochastic method) shows a rapid initial increase on a timescale $t\sim\varepsilon$, followed by a slower approach to the final equilibrium value. The high-temperature (white-noise) prediction lacks this initial rapid rise entirely, as it misses the dominant contribution from high-frequency bath modes ($\omega\sim\varepsilon^{-1}$).
This behavior can be understood analytically from the low-temperature limit of the momentum variance breuer :
$\langle p^{2}(t)\rangle\approx\frac{m\gamma\hbar}{\pi}\int_{0}^{\infty}d\omega\,\omega e^{-\varepsilon\omega}\frac{|e^{i\omega t}-e^{-\gamma t}|^{2}}{\omega^{2}+\gamma^{2}}.$ (45)
For short times $t\ll\varepsilon$, the integral yields $\langle p^{2}(t)\rangle-\langle p^{2}(0)\rangle\propto(m\gamma\hbar/\pi)(t^{2}/\varepsilon^{2})$. For intermediate times $\varepsilon\ll t\ll\gamma^{-1}$, the growth is logarithmic: $\propto(2m\gamma\hbar/\pi)(1-\gamma t)\ln(t/\varepsilon)$. Thus, the bath cutoff $\varepsilon^{-1}$ sets the timescale for the initial, rapid thermalization driven by vacuum fluctuations—a genuinely quantum effect that ensures fast decoherence even at $T=0$.
Figure 3: Momentum dispersion $\langle p^{2}(t)\rangle$ relaxation for Schrödinger cat state. The exact quantum result (top lines, analytical and stochastic) shows rapid initial growth on a timescale $\sim\varepsilon$, absent in the high-temperature white-noise model (dashed). This demonstrates the quantum bath’s role in fast early-time dynamics. Parameters: $\gamma=\pi/2$, $\varepsilon=0.01$, $T=1$, $m=\hbar=1$.
VI Extension to Non-Quadratic Potentials: A Controlled Approximation
The exact quantum-classical correspondence demonstrated in the previous sections holds strictly for Hamiltonians that are quadratic in both the particle and bath coordinates. For an arbitrary external potential $V(x)$, this exact correspondence is broken. Remarkably, however, the intrinsic structure of the Caldeira-Leggett model provides a natural mechanism that suppresses quantum coherence, rendering the dynamics increasingly classical and thus validating the stochastic approach as a controlled approximation.
VI.1 A Natural Small Parameter: The Equilibrium Coherence Length
The key lies in analyzing the equilibrium reduced density matrix of the particle, $\rho_{S}^{\text{eq}}(x,x^{\prime})=\rho_{S}^{\text{eq}}(r,q)$, derived from the full thermal state $\rho_{SB}^{\beta}$. For the CL model, this density matrix develops an exponential decay in its off-diagonal coordinate $q=x-x^{\prime}$ Weiss_2012 :
$\rho_{S}^{\text{eq}}(r,q)\propto\exp\!\left(-\frac{1}{2\hbar^{2}}\langle p^{2}\rangle_{\text{eq}}\,q^{2}\right),$ (46)
where $\langle p^{2}\rangle_{\text{eq}}$ is the equilibrium momentum variance of the particle, given by (45) in the limit of a wide bath spectrum. This defines a coherence length $\lambda$:
$\lambda\equiv\frac{\hbar}{\sqrt{\langle p^{2}\rangle_{\text{eq}}}}.$ (47)
The physical significance of $\lambda$ is clear from Eq. (46): the reduced density matrix becomes strongly peaked along the diagonal $q=0$ (i.e., $x\approx x^{\prime}$) when $\lambda$ is small. The off-diagonal elements, which encode quantum coherence, are exponentially suppressed for $|q|\gg\lambda$.
Crucially, for an ohmic bath with a high-frequency cutoff $\Lambda\sim\varepsilon^{-1}$, the momentum variance diverges in the wide-band limit:
$\langle p^{2}\rangle_{\text{eq}}\propto\int_{0}^{\Lambda}d\omega\,\frac{J(\omega)}{\omega}\coth\left(\frac{\hbar\omega}{2k_{B}T}\right)\xrightarrow[\Lambda\to\infty$$\infty$. (48)
Consequently, the coherence length vanishes:
$\lambda\xrightarrow[\Lambda\to\infty]{}0.$ (49)
This occurs even at zero temperature due to coupling to vacuum fluctuations, unlike the thermal coherence length $\lambda_{\text{th}}\sim\hbar/\sqrt{mk_{B}T}$, which remains finite as $T\to 0$. Thus, $\lambda$ emerges as a natural, bath-controlled small parameter that quantifies the degree of ”classicality” induced in the particle by its interaction with a broad-spectrum environment.
VI.2 Validity of the Stochastic Approximation
The stochastic approach derived for quadratic potentials can be extended to a smooth, non-quadratic potential $V(x)$ when the system operates in the quasi-classical regime. This regime is defined by the condition that the potential varies slowly over the scale of the quantum coherence length $\lambda$.
Let $L$ be the characteristic classical length scale of the potential, which can be defined via its derivatives, e.g., $L\sim|V^{\prime}(x)/V^{\prime\prime\prime}(x)|^{1/2}$ or the scale over which $V(x)$ changes appreciably. The central result is that the error incurred by approximating the exact quantum dynamics with the classical stochastic process is controlled by the dimensionless ratio
$\epsilon\equiv\frac{\lambda}{L}\ll 1.$ (50)
When $\epsilon\ll 1$, the particle’s density matrix remains nearly diagonal throughout its evolution, as any generated coherences are rapidly suppressed by the bath on the scale $\lambda$. In this regime, the dominant contributions to path integrals or master equations come from nearly diagonal histories, which are accurately captured by the ensemble of classical trajectories governed by the GLE with quantum noise.
Therefore, for potentials satisfying Eq. (50), we expect the stochastic method to provide an excellent approximation to the true quantum dynamics, with deviations of order $\mathcal{O}(\epsilon)$. This generalizes the utility of our approach beyond exactly solvable models, providing a powerful and practical tool for studying dissipative quantum dynamics in complex potentials, from chemical reaction profiles to disordered systems.
VI.3 Outlook and Future Development
This analysis provides a clear pathway for future work. The next step is to develop a systematic perturbation theory in the small parameter $\epsilon=\lambda/L$. This would allow for the calculation of leading quantum corrections to the classical stochastic dynamics, potentially through the inclusion of non-local (in time) correction terms to the stochastic weights or via a modified Fokker-Planck equation. Such a framework would bridge the gap between the exact correspondence established for quadratic systems and the approximate but highly efficient stochastic simulation available for a much broader class of problems.
VII Conclusions
The central challenge in simulating quantum Brownian motion (QBM) is the exact treatment of environmental influence, which makes direct quantum mechanical calculations intractable due to the exponentially large Hilbert space. While the path-integral framework provides an exact solution for the quadratic Caldeira-Leggett (CL) model, the resulting expressions are cumbersome and resist extension to anharmonic potentials or complex preparation protocols.
In this work, we have introduced a profound reinterpretation and a powerful numerical method that overcomes these barriers. We have demonstrated that the exact quantum dynamics of a particle in the CL model can be mapped, at any temperature, onto a classical, non-Markovian stochastic process in phase space. For a particle with a quadratic Hamiltonian, this correspondence is proven exact, starting from the physically correct correlated thermal equilibrium state and accommodating arbitrary state preparations and interventions via the Wigner representation of preparation functions.
The primary value of this mapping is its profound simplification: it reduces the formidable problem of simulating an open quantum system to the tractable task of generating and averaging classical stochastic trajectories. We have developed and validated a corresponding Monte Carlo numerical technique. The method correctly reproduces established analytical results and, crucially, captures the fast decoherence and thermalization dynamics driven by vacuum and low-temperature fluctuations—effects that are missed by common high-temperature master equations.
Furthermore, we have identified a natural small parameter that governs the extension of this framework: the equilibrium coherence length $\lambda=\hbar/\sqrt{\langle p^{2}\rangle_{\text{eq}}}$, which shrinks with increasing bath spectral width. This analysis shows that for smooth, non-quadratic potentials where the classical scale $L$ satisfies $\lambda/L\ll 1$, our stochastic approach provides a controlled approximation, with error of order $\mathcal{O}(\lambda/L)$. This opens a clear pathway for future development, including systematic perturbation theory or hybrid numerical schemes that treat confined quantum dynamics exactly within the coherence width $\lambda$ while sampling the larger-scale classical motion stochastically.
In summary, we have established a versatile, all-temperature computational framework for QBM. By translating non-Markovian quantum dynamics into a classical stochastic language, we provide a conceptually simpler and numerically efficient alternative to more complex contemporary methods, offering a unified tool for studying dissipative quantum dynamics from the classical high-temperature regime down to the deeply quantum zero-temperature limit.
Acknowledgement
This work was supported by Rosatom in the framework of the Roadmap for Quantum computing (Contract No. 868-1.3-15/15-2021 dated October 5).
Appendix A Influence functional
A.1 Classic non-Markovian stochastic process
For generalized Langevin equation Hanggi1993
$m\ddot{x}+\int_{t_{i}}^{t}dsM(t-s)\dot{x}(s)=\xi(t)$ (51)
with colored noise $\langle\xi(s)\xi(s+t)\rangle=\sigma(t)$ propagator (with the normalized prefactor omitted) is written as
$K(x_{f}p_{f}t_{f}|x_{i}p_{i}t_{i})=\int_{x_{i}}^{x_{f}}\mathcal{D}x\int\mathcal{D}z\delta(\dot{x}_{i}-\frac{p_{i}}{m})\delta(\dot{x}_{f}-\frac{p_{f}}{m})e^{iS[x,z]},$ (52)
where effective action is
$\begin{gathered}S[x,z]=\int_{t_{i}}^{t_{f}}dsz(s)[\ddot{x}(s)+\int_{t_{i}}^{s}du\gamma(s-u)\dot{x}(u)]+\\ +\frac{i}{2}\int_{t_{i}}^{t_{f}}ds\int_{t_{i}}^{t_{f}}duz(s)\sigma(s-u)z(u).\end{gathered}$ (53)
with $\gamma(t)=\frac{1}{m}M(t)$
A.2 Quantum influence functional with factorized initial condition
For non-correlated state we have
$\rho_{0}^{SB}=\rho_{0}\otimes\rho^{B}_{\beta}$ (54)
$J(r_{f}q_{f}t_{f}|r_{i}q_{i}t_{i})=\int_{r_{i}}^{r_{f}}Dr\int_{q_{i}}^{q_{f}}Dqe^{\frac{i}{\hbar}A[r,q]}$ (55)
$\begin{gathered}A[r,q]=-\int_{t_{i}}^{t_{f}}dsmq(s)[\ddot{r}(s)+\int_{t_{i}}^{s}du\gamma(s-u)\dot{r}(u)]+\\ +m(\dot{r}_{f}q_{f}-\dot{r}_{i}q_{i})+\frac{i}{2\hbar}\int_{t_{i}}^{t_{f}}ds\int_{t_{i}}^{t_{f}}duq(s)N(s-u)q(u)\end{gathered}$ (56)
Go to Wigner representation $W(rpt)=\frac{1}{2\pi\hbar}\int dq\rho(rqt)e^{-\frac{i}{\hbar}pq}$ in
$\rho(r_{f}q_{f}t-F)=\int dr_{i}dq_{i}J(r_{f}q_{f}t_{f}|r_{i}q_{i}t_{i})\rho(r_{i}q_{i}t_{i})$ (57)
- multiply by $e^{-ipq/\hbar}$ and integrate by $q$, also represent initial density matrix through Wigner function
$W(r_{f}p_{f}t_{f})=\int{dq_{f}dr_{i}dq_{i}dp_{i}J(r_{f}q_{f}t_{f}|r_{i}q_{i}t_{i})}\\ \times e^{-\frac{i}{\hbar}(p_{f}q_{f}-p_{i}q_{i})}W(r_{i}p_{i}t_{i})$ (58)
So we obtain Wigner representation of propagator
$K_{Q}(r_{f}p_{f}t_{f}|r_{i}p_{i}t_{i})=\int dq_{i}dq_{f}J(r_{f}q_{f}t_{f}|r_{i}q_{i}t_{i})e^{-\frac{i}{\hbar}(p_{f}q_{f}-p_{i}q_{i})}$ (59)
Finally we have
$K_{C}(x_{f}p_{f}t_{f}|x_{i}p_{i}t_{i})=\int_{x_{i}}^{x_{f}}Dx\int Dz\delta(\dot{x}_{i}-\frac{p_{i}}{m})\delta(\dot{x}_{f}-\frac{p_{f}}{m})\\ \times\exp\left[i\int_{t_{i}}^{t_{f}}dsz(s)\left(\ddot{x}(s)+\int_{t_{i}}^{s}du\gamma(s-u)\dot{x}(u)\right)\right.\\ \left.-\frac{1}{2}\int_{t_{i}}^{t_{f}}ds\int_{t_{i}}^{t_{f}}duz(s)\sigma(s-u)z(u)\right]$ (60)
- propagator, constructed for initially classical process
$K_{Q}(r_{f}p_{f}t_{f}|r_{i}p_{i}t_{i})=\int dq_{i}dq_{f}e^{-\frac{i}{\hbar}(p_{f}q_{f}-p_{i}q_{i})}\int_{r_{i}}^{r_{f}}Dr\int_{q_{i}}^{q_{f}}Dq\\ \exp[-\frac{im}{\hbar}\int_{t_{i}}^{t_{f}}dsq(s)\bigg(\ddot{r}(s)+\int_{t_{i}}^{s}du\gamma(s-u)\dot{r}(u)\bigg)\\ -\frac{1}{2\hbar^{2}}\int_{t_{i}}^{t_{f}}ds\int_{t_{i}}^{t_{f}}duq(s)N(s-u)q(u)\bigg]e^{\frac{im}{\hbar}(\dot{r}_{f}q_{f}-\dot{r}_{i}q_{i})}$ (61)
- propagator, obtained by Wigner transform from exactly quantum dynamics.
Calculate explicitly $\int Dq$, slicing $[0,t]$ by intervals $\tau=t/N:[t_{0},t_{1},...,t_{N}]$, so $q(t_{k})=q_{k}$. Measure
$Dq=\lim_{N\to\infty}\prod_{k=1}^{N}\frac{dq_{k}}{A},$ (62)
where $A$ - normilized factor and boundary conditions $q_{0}=q_{i},q_{N}=q_{f}$
Discrete integral by DqDq
$\int\prod_{k=1}^{N-1}\frac{dq_{k}}{A}\exp[-\frac{im}{\hbar}\sum_{k=0}^{N}\tau q_{k}\bigg(\ddot{r}_{k}+\sum_{j=0}^{k}\gamma_{kj}\dot{r}_{k}\bigg)\\ -\frac{1}{2\hbar^{2}}\sum_{k=0}^{N}\sum_{j=0}^{N}\tau^{2}q_{k}N_{kj}q_{j}\bigg],$ (63)
where $\gamma_{kj}=\gamma(t_{k}-t_{j})$, $N_{kj}=N(t_{k}-t_{j})$. Let us select in the sums the parts containing the boundary terms $q_{i},\,q_{f}$
$\int\prod_{k=1}^{N-1}\frac{dq_{k}}{A}\exp[-\frac{im\tau}{\hbar}q_{0}[\ddot{r}_{0}+\sum\gamma_{0j}\dot{r}_{0}]\\ -\frac{im\tau}{\hbar}q_{N}[\ddot{r}_{N}+\sum\gamma_{Nj}\dot{r}_{N}]\\ -\frac{im\tau}{\hbar}\sum_{k=1}^{N-1}q_{k}[\ddot{r}_{k}+\sum\gamma_{kj}\dot{r}_{k}]-\frac{\tau^{2}}{2\hbar^{2}}\sum_{k=1}^{N-1}\sum_{j=1}^{N-1}q_{k}N_{kj}q_{j}\\ -\frac{\tau^{2}}{2\hbar^{2}}q_{0}N_{00}q_{0}-\frac{\tau^{2}}{2\hbar^{2}}q_{N}N_{NN}q_{N}-\frac{\tau^{2}}{2\hbar^{2}}\sum_{j=1}^{N-1}q_{0}N_{0j}q_{j}\\ -\frac{\tau^{2}}{2\hbar^{2}}\sum_{k=1}^{N-1}q_{k}N_{k0}q_{0}-\frac{\tau^{2}}{2\hbar^{2}}\sum_{j=1}^{N-1}q_{N}N_{Nj}q_{j}-\frac{\tau^{2}}{2\hbar^{2}}\sum_{k=1}^{N-1}q_{k}N_{kN}q_{N}\bigg]$ (64)
We use the symmetry of $N_{kj}$ and also note that the terms containing only $q_{i}$ and $q_{f}$ are multiplied by $\tau$ and vanish in the limit. The difference arises in the sums, where the boundary terms are linked with the integrable variables. We combine similar terms
$\int\prod_{k=1}^{N-1}\frac{dq_{k}}{A}\exp[-\frac{im\tau}{\hbar}\sum_{k=1}^{N-1}q_{k}\Big([\ddot{r}_{k}+\sum\gamma_{kj}\dot{r}_{k}]\\ -\frac{i\tau}{\hbar m}[N_{k0}q_{0}+N_{kN}q_{N}]\Big)-\frac{\tau^{2}}{2\hbar^{2}}\sum_{k=1}^{N-1}\sum_{j=1}^{N-1}q_{k}N_{kj}q_{j}\bigg]$ (65)
The term dependent on the boundary coordinates disappears in the limit $\tau\xrightarrow[]{}0$, so the boundary conditions $q_{i}$ and $q_{f}$ are not included in the last integral and the dependence on them remains only in the first two exponentials, the integration of which yields delta functions $\delta(p-m\dot{r}_{f})\delta(p_{0}-m\dot{r}_{i})$. So with precision by variables replacing, $K_{Q}=K_{C}$. Because we obtain the identity in path integral form, we can also reproduce any multitime correlator functions, so we proove full equality of classic and quantum processes in this case.
A.3 Replacing integration in imaginary time with a negative time for correlated initial states
To comparisong with Grabert1988 we show, as it taken into account correlation between particle and environment in $t=0$. Notice that $\rho_{\beta}^{SB}$ is stationary state of full Hamiltionian, so we can identically rewrite it as
$U(0,-T)\rho_{\beta}^{SB}U(-T,0)$ (66)
Because of preparation acts only to particle, trajectories of oscillators remains continuous and after tracing by them includes into effective action analogously, but with history from T-T to $t$, and trajectory of particle has a jump in $t=0$. Directing $T\xrightarrow$$\infty$, terms connecting dynamics with initial state in t=Tt=-T are dissapeared, that allows consider arbitrary state as initial. Our numerical approach based on this trick.